# An End-to-End Homomorphically Encrypted Neural Network

### v1

@@ Abstract
*Equal contribution. Listing order is random. Florencio proposed using homomorphic encryption on neural networks and was involved in every aspect of this work. He also designed the first version of the Differentiable Soft-Argmax layer. Alencar implemented the final versions of the training pipelines and the Differentiable Soft-Argmax layer, and was responsible for the experiments. Lima was in charge of putting together the final version of the paper and the experiments and also spent countless hours developing an interactive demo for the model, available at https://www.epiphyte-2.com.
@@ 1 Introduction
At the same time that neural networks are considered ”black-boxes,” due to their inherent internal complexity, the need of relinquishing privacy in order to use state-of-the-art neural networks is a well-known issue, especially for companies that deal with sensitive data [4, 17]. In these models, the input data is processed in plain text and then converted into a vector of features that is used by the model to perform inference. If the input could be encrypted, the model could be used without the need of sharing the plain data with the model owner. In an end-to-end homomorphically encrypted neural network, the input data is encrypted and the model operates on the encrypted data, providing full privacy and security.
@@
This result can be achieved by adapting the encryption scheme proposed by Cheon et al. (2016), known as HEANN or CKKS, which is designed to perform computation on real numbers, which happens to be the format of the input and output of neural networks. The main idea is to encrypt the input data, plus some acceptable level of noise, and then calibrate the model’s weights in such a way that the forward pass of the neural network can be performed in the encrypted domain. In the same fashion, the outputs of the model are calibrated so that only the owner of the private key can decrypt the output. The main challenge of this approach is the noise that is introduced in the encrypted domain, which can be mitigated by using a noise budget, which is a parameter that limits the amount of noise that can be introduced in the encrypted domain. Fortunately, while the base scheme requires specific Add and Mult algorithms to perform operations, in the context of neural networks they are performed internally and the noise budget can be controlled by the model’s architecture and the training process during backpropagation.
@@
Basically, in all neural network architectures, the output of the model is a vector of logits, which are the raw outputs of the model before the softmax function. These logits are used to calculate the loss function and to perform inference. It is our understanding that a privacy-preserving neural network should maintain the output logits in the encrypted domain, ideally introducing new noise, in such a way that the logits can still be used to infer information about the input data. To mitigate this issue, we propose a new layer, the Differentiable Soft-Argmax, which is a differentiable version of the argmax function, that can be used to calibrate the logits in the encrypted domain, raising the entropy of the activation parameters, thus improving the security of the model while keeping the overall noise below the acceptable noise budget.
@@
As best as we can determine, we believe that this is the first work that proposes an end-to-end homomorphically encrypted neural network relying solely on encryption and backpropagation to train the model. We also believe that this is the first work that formally introduces a new layer, which we call Differentiable Soft-Argmax, that can be used to calibrate the output logits in the encrypted domain.
@@ 3.2 Modified HEANN
Also as suggested by Gentry [8, p. 6], we use the parameter $B$ to determine the formatting of the secret key and ciphertexts as inputs and outputs of the network, whose size is a fixed polynomial in the security parameter, meaning the ciphertext size depends only on the security parameter and is independent of the circuit $C$.
@@ 7 Conclusion
In this work, we presented what we believe to be the first functioning homomorphically encrypted neural network that can operate on encrypted data while maintaining high levels of accuracy. We also introduced a new layer, the Differentiable Soft Argmax, which allows the calibration of output logits in the encrypted domain, raising the entropy of the activation parameters and improving the security of the model. In his seminal thesis (p.8), Gentry highlights that constructing an efficient homomorphic encryption scheme without using bootstrapping, or using some relaxations of it, is an interesting open problem (p.9). At least in the context of neural networks, that is what we believe to have accomplished.
For sentiment analysis tasks, our experiments show that the HNN model can achieve up to 82.5% of the accuracy of the plain model while maintaining full privacy and security. We believe that this work represents a significant step forward in the field of privacy-preserving neural networks and that it has the potential to open up new avenues for research in this area.
We are excited about the potential applications of this technology and look forward to further exploring its capabilities in the future. We plan to continue our research by investigating how mechanistic interpretability can be achieved in the context of homomorphically encrypted neural networks and how safe multi-party computation can be integrated into our models and pipelines.

### v2

@@ Abstract
This is an update of the version published in February 22, 2025, due to the inclusion of patent number BR 10 2025 003497 2.
*Equal contribution. Listing order is random. Florencio proposed the use of homomorphic encryption in neural networks and was involved in all aspects of this work. He also designed the first version of the Differentiable Soft-Argmax layer. Alencar implemented the final versions of the training pipelines and the Differentiable Soft-Argmax layer, and was responsible for the experiments. Lima was in charge of putting together the final version of the paper and the experiments and also spent countless hours developing an interactive demo for the model, available at https://www.epiphyte-2.com.
@@ 1 Introduction
At the same time that neural networks are considered ”black-boxes,” due to their inherent internal complexity, the need to relinquish privacy to use state-of-the-art neural networks is a well-known issue, especially for companies that deal with sensitive data [5, 24]. In these models, the input of plain text is converted into a vector of features that the model uses to perform inference. If the input could be encrypted, the model could be used without the need to share the plain data with the model owner, providing full privacy and security.
@@
We adapt the encryption scheme proposed by [6], which is designed to perform computation on real numbers, to encrypt input data, plus some acceptable level of noise, and then calibrate the model’s weights in such a way that the forward pass of the neural network can be performed in the encrypted domain. In the same fashion, the outputs of the model are calibrated so that only the owner of the private key can decrypt the output. The main challenge of this approach is the noise that is introduced in the encrypted domain, which can be mitigated by using a noise budget, which is a parameter that limits the maximum amount of noise. While the base scheme requires specific Add and Mult algorithms to perform operations, they are performed internally in the context of neural networks and the noise budget is controlled by the model’s architecture and the training process during backpropagation.
@@
It is also commom practice in neural network architectures that the output of the model is a vector of raw logits, upon which the softmax function is applied to calculate the loss function and to perform inference[13, 20, 28, 26]. It is our understanding that a privacy-preserving neural network should maintain the output logits in the encrypted domain, ideally introducing new noise. To address this issue, we propose a new layer, the Differentiable Soft-Argmax, which is a differentiable version of the argmax function. It calibrates the logits in the encrypted domain, raising the entropy of the activation parameters, thus improving the security of the model while keeping the overall noise below the acceptable noise budget.
2 Background
Homomorphically encrypted neural networks is a relatively new field of research, with the first significant work being published in 2016 [11]. It demonstrated how a network can execute over encrypted data and introduced optimizations to make inference feasible. The encryption scheme chosen was YASHE’, described in [2], which does not support floating-point numbers and allows operations on encrypted data if, and only if, the complexity of the arithmetic circuit is known in advance.
Both of these limitations were addressed by [6], which introduced a new encryption scheme that allows operations on encrypted data without the need of knowing the complexity of the arithmetic circuit in advance. This scheme is known as HEANN or CKKS, and it is the one we use in this work. Another fundamental difference is the optizations proposed by [11], such as replacing activation functions by polynomials, which end up tampering with the model’s architecture, making their approach less generalizable. We rely solely on backpropagation to perform all operations internally while keeping the noise budget controlled. This is acceptable because, even though the weights are known, the underlying embeddings are unknown, due to the noise added during encryption.
Since then, several works have been published, such as [15], which implements word-wise encryption and replaces activation functions by precise polynomial approximation techniques, [16], which proposes quantization as a way to allow more efficient computations, and [19], which focuses on using homomorphic encryption as a solution to build training approachs on encrypted data combined with multi-party computation.
@@ 2 Background
As best we can determine, we believe that this is the first work that proposes an end-to-end homomorphically encrypted neural network that relies solely on encryption and backpropagation to train the model. We also believe that this is the first work that formally introduces a new layer, which we call Differentiable Soft-Argmax, that can be used to calibrate the output logits in the encrypted domain.
@@ 4.2 Modified HEANN
Also as suggested by Gentry [9, p. 6], we use the parameter $B$ to determine the formatting of the secret key and ciphertexts as inputs and outputs of the network, whose size is a fixed polynomial in the security parameter, meaning the ciphertext size depends only on the security parameter and is independent of the circuit $C$.
@@ 8 Conclusion
In this work, we presented a functioning homomorphically encrypted neural network that can operate on encrypted data while maintaining high levels of accuracy. We also introduced a new layer, the Differentiable Soft Argmax, which allows the calibration of output logits in the encrypted domain, raising the entropy of the activation parameters and improving the security of the model. In his seminal thesis, Gentry highlights that constructing an efficient homomorphic encryption scheme without using bootstrapping, or using some relaxations of it, is an interesting open problem (p.9). At least in the context of neural networks, that is what we believe to have accomplished.
For sentiment analysis tasks, our experiments show that the HNN model can achieve up to 82.5% of the accuracy of the plain model while maintaining full privacy and security. We believe that this level of accuracy, while significant, can be improved by further refining the training pipelines and methodologies.
We are excited about the potential applications of this technology and look forward to further exploring its capabilities. We plan to continue our research by investigating how mechanistic interpretability can be achieved in the context of homomorphically encrypted neural networks and how safe multi-party computation can be integrated into our models and pipelines.